### v1

# Characterization of Efficient Influence Function for Off-Policy Evaluation Under Optimal Policies

Abstract
Off-policy evaluation (OPE) provides a powerful framework for estimating the value of a counterfactual policy using observational data, without the need for additional experimentation. Despite recent progress in robust and efficient OPE across various settings, rigorous efficiency analysis of OPE under an estimated optimal policy remains limited. In this paper, we establish a concise characterization of the efficient influence function for the value function under optimal policy within canonical Markov decision process models. Specifically, we provide the sufficient conditions for the existence of the efficient influence function and characterize its expression. We also give the conditions under which the EIF does not exist.
keywords:
Semiparametric Efficiency; Efficient Influence Function; Optimal Policy; Off-Policy Evaluation.
journal: arxiv
1 Introduction
Reinforcement learning (RL) focusing on developing optimal policies for sequential decision-making to maximize long-term rewards, (Sutton & Barto, 2018) have been serving as more and more important frontier in various fields. A critical component of RL is off-policy evaluation (OPE), which estimates the mean reward of a policy, termed the evaluation policy, using data collected under another policy, known as the behavior policy. OPE is essential in offline RL, where only historical datasets are available, precluding new experiments (Luedtke & Van Der Laan, 2016; Agarwal et al., 2019; Uehara et al., 2022).
Recent years have witnessed substantial progress in developing statistically efficient OPE methods, with various approaches demonstrating semiparametric efficiency under different model settings (Jiang & Li, 2016; Kallus & Uehara, 2020; Shi et al., 2021). However, all of these existing analyses focus on scenarios where the evaluation policy is fixed and predetermined. A more challenging yet practical scenario arises when the evaluation policy itself is estimated from data, particularly when this policy is designed to be optimal with respect to some criterion. In this context, the statistical properties of OPE become more complex due to the additional estimation uncertainty introduced by the policy optimization process.
In contrast, in the causal inference literature, such phenomena have been studied extensively in the optimal treatment regime literature (Laber et al., 2014; Kosorok & Laber, 2019; Athey & Wager, 2021). These works have established important results regarding the estimation of value functions under optimal treatment rules, but their direct application to the sequential decision-making context of RL presents additional challenges due to the temporal dependencies and potentially infinite horizons involved.
In this paper, we provide a simple result for the characterization of the efficient influence function (EIF) for the value function under optimal policy within canonical Markov decision process models. Specifically, we establish sufficient conditions for the existence of the EIF and derive its explicit expression. We also identify and prove conditions under which the EIF does not exist, and demonstrate that various existing methods, including the value functions in Fitted Q Iteration (FQI, for example Munos & Szepesvári, 2008; Shi et al., 2024a) and Fitted Actor-Critic (FAC, see Xu et al., 2021; Kumar et al., 2023), do achieve semiparametric efficiency without parameterization or fixing the policies.
2 Problem Formulation
2.1 Data Generating Process and Parameter of Interest
We consider observational data generated from a canonical Markov decision process (MDP). Specifically, at a given time $t$, let $(S_{t},A_{t},R_{t})$ denote the observed state-action-reward tripleruplet. Let $O$ be a shorthand for a data tuple $(S,A,R,S^{\prime})$. Given an offline observable dataset $\big{\{}O_{it}:1\leq i\leq N,0\leq t\leq T\big{\}}$ with $O_{it}=(S_{it},A_{it},R_{it})$ generated by a behavior policy $b(\cdot\mid S)$, where $i$ indexes the $i$-th episode and $t$ indexes the $t$-th time point. For any fixed target policy $\pi(a\mid s)$, OPE geneally aim to evaluate the mean return under this fixed policy as
$\eta(\pi)=\mathrm{E}^{\sim\pi}\bigg{[}\sum_{t=0}^{+\infty}\gamma^{t}R_{t}\bigg{]}$
with a valid confidence interval, where $\mathrm{E}^{\sim\pi}$ denotes expectation when the system follows the policy $\pi$. Slightly different from the existent semiparametric study on OPE, instead of considering the arbitrary $\pi$, in this paper we will focus on a specific target policy: the optimal policy $\pi^{*}$, which maximize the $\eta(\pi)$ in the set of all possible policies $\mathcal{P}$. Specially, the parameter of insterests we will focus on is
$\eta^{*}:=\eta(\pi^{*})=\mathrm{E}^{\sim\pi^{*}}\bigg{[}\sum_{t=0}^{+\infty}\gamma^{t}R_{t}\bigg{]}\quad\text{ such that }\quad\pi^{*}=\arg\max_{\pi\in\mathcal{P}}\eta(\pi).$
To ensure the value function is identifiable, as commonly used in various OPE literature, we present the following assumptions. For simplicity, we use $f(x\mid y)$ to represent the conditional density of $X$ given $Y=y$.
{assumption}
[Data Structure & Observations] The observations are i.i.d. copies of the trajectory $\{(S_{t},A_{t},R_{t},S_{t+1})\}_{t\geq 0}$, following the data-generating process (DGP):
$S_{t+1}\sim f(s_{t+1}\mid A_{t},S_{t}),\qquad R_{t}\sim f(r_{t}\mid A_{t},S_{t}),\qquad A_{t}\sim b(a_{t}\mid S_{t}).$
{assumption}
[Markov, Conditional Independence, & Time-Homogeneity] $f(a_{t},s_{t}\mid a_{t-1},s_{t-1},a_{t-2},s_{t-2},\ldots)=f(a_{t},s_{t}\mid a_{t-1},s_{t-1})$ for any $t\geq 1$; $f(a_{t}\mid a_{t-1},s_{t})=b(a_{t}\mid s_{t})$. The reward $R_{t}$ only depends on $A_{t}$ and $S_{t}$; All conditional density functions $b(a\mid s)$, $f(r\mid a,s)$, and $f(s^{\prime}\mid a,s)$ remain fixed over time.
2.2 Two Estimators under Fixed Policies
As shown in various literature, Assumption 2.1 and Assumption 2.1 is sufficient for identifying $\eta(\pi)$ for any given policy $\pi\in\mathcal{P}$. Here we first discuss the commonly-used estimating methods.
The first estimating method involves analyzing the aggregate mean return under any target policy $\pi$, and we introduce two “state-action” value functions, commonly referred to as the $Q$-function. The $Q$-function evaluates the expected cumulative reward given a specific action at a specific state (Uehara et al., 2022) under a fixed hypothetical policy $\pi$, defined as
$Q(a,s;\pi):=\mathrm{E}^{\sim\pi}\bigg{[}\sum_{k=0}^{+\infty}\gamma^{k}R_{t+k}\mid A_{t}=a,S_{t}=s\bigg{]}$ (1)
Then the value function can be rewritten as
$\eta(\pi)=\mathrm{E}^{\sim\pi}\Bigg{[}\mathrm{E}^{\sim\pi}\bigg{[}\sum_{t=0}^{+\infty}\gamma^{t}R_{t}\mid A_{0},S_{0}\bigg{]}\Bigg{]}=\int Q(a_{0},s_{0};\pi)\pi(a_{0}\mid s_{0})f(s_{0})\mathrm{d}a_{0}\,\mathrm{d}s_{0},$
under any policy $\pi$ where $f(s_{0})$ is the density of the initial state distribution. Usually, (LABEL:Q_to_eta_eq) ensures it is sufficient to learn the cumulative rewards for any given policy of interest $\pi$, and we also define a weighted average of the $Q$-function as $V(s;\pi)=\int Q(a,s;\pi)\pi(a\mid s)\,\mathrm{d}a$.
The second estimator is the marginal importance sampling (MIS) estimator, which aims to address the curse of horizon. The marginal ratio is defined as
$\omega(s;\pi):=(1-\gamma)\sum_{t=0}^{+\infty}\frac{\gamma^{t}f_{\sim\pi,t}(s)}{f_{+\infty}(s)},$ (2)
where $f_{\sim\pi,t}(s)$ is the density function of the state $S_{t}$ under policy $\pi$, and $f_{+\infty}(s)$ is the marginal density function of the stationary state distribution $S_{+\infty}$. Under the stationarity assumption, we have the following identity
$\eta(\pi)=\frac{1}{1-\gamma}\mathrm{E}\big{[}\omega(S;\pi)\mathrm{E}^{\sim\pi}[R\mid S]\big{]}.$
An important observation is that both the $Q$-function (1) and MIS ratio (2) are included in the semiparametric efficient estimator for $\eta(\pi)$.
2.3 Characterization and Issues
Assume the trajectory $O=O_{1:T}\sim P_{0}\in\mathcal{M}$ and let $\Psi^{*}:\mathcal{M}\rightarrow\mathbb{R}$ be defined as
$\Psi^{*}(P):=\mathrm{E}_{P}\big{[}Q(P)\big{(}A_{0},S_{0};\pi^{*}(P)\big{)}\pi^{*}(P)(A_{0}\mid S_{0})\big{]}$
where $\pi^{*}(P)(\cdot\mid s)=\arg\max_{\pi\in\mathcal{P}}Q(P)(a,s;\pi)$ is the optimal policy under the law $P$. Then $\Psi^{*}(P_{0})=\eta(\pi^{*})$ is well-defined. Instead of focusing on $\pi^{*}(P_{0})$ itself, here we only focus on the value function under the optimal policy $\Psi^{*}(P_{0})$, whereas the discussion for $\pi^{*}(P_{0})$ can be found in Kosorok & Laber (2019); Athey & Wager (2021); Luo et al. (2024). Define the suboptimal gap between $\pi$ and the optimal policy $\pi^{*}:=\pi^{*}(P)$ as
$\Delta(P)(\pi;A,S):=Q(P)(A,S;\pi^{*})-Q(P)(A,S;\pi)$
which implies $\pi^{*}(A\mid S):=\arg\min_{\pi\in\mathcal{P}}\Delta(\pi;A,S)=\pi^{*}(P)(A,S)$. Define the new functional
$\Psi(P;\pi)=\mathrm{E}_{P}\big{[}Q(P)\big{(}A_{0},S_{0};\pi\big{)}\pi(A_{0}\mid S_{0})\big{]}$
then $\Psi\big{(}P;\pi^{*}(P)\big{)}=\Psi^{*}(P)$, but in general, we will not have $\Psi\big{(}P_{1};\pi^{*}(P_{2})\big{)}=\Psi^{*}(P_{1})$ if $P_{1}\neq P_{2}$.
Suppose $\{S_{t}\}_{t\geq 0}$ is stationary, as shown in Uehara et al. (2020); Shi et al. (2024b, 2021), when for any fixed $\pi$, the efficient influence function (EIF) for $\Psi(P;\pi)$ at $P_{0}$ evaluated with $O$ on the full non-parametric space $\mathcal{M}_{\text{nonpar}}$ is
$S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\Big{|}_{P=P_{0}}(O)$ (3)
$=$ $\frac{1}{1-\gamma}\omega(P_{0})(S;\pi)\frac{\pi(A\mid S)}{b(P_{0})(A\mid S)}\big{[}R+\gamma V(P_{0})(S^{\prime};\pi)-Q(P_{0})(A,S;\pi)\big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+V(P_{0})(S;\pi)-\Psi(P_{0};\pi).$
Since for any regular asymptotic normal (RAL) estimator $\widehat{\Psi}(\widehat{P};\pi)$, there exists a unqiue influence function (IF, Tsiatis, 2006) such that
$\widehat{\Psi}(\widehat{P};\pi)-\Psi(P_{0};\pi)=\sqrt{N}\mathbb{P}_{N}\operatorname{IF}(P_{0};\pi)(O)+o_{P_{0}}\big{(}N^{-1/2}\big{)},$
the EIF $S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0}}(O)$ serves an crucial role in semiparametric statistics, as it is a vaid IF minimizes the variance $\operatorname{var}_{P_{0}}\big{(}\operatorname{IF}(P_{0};\pi)(O)\big{)}$.
The set of all valid influence functions $\operatorname{IF}(P_{0};\pi)(O)$ is uniquely characterized by the tangent space $\mathcal{T}$, i.e., the mean-square closure of the parametric submodels (for both nuisance tangent space and full tangent space), which is equivalently characterized by the collection of score functions of any differentiable paths at $P_{0}$ (see discussion in Van Der Vaart et al., 1996). Thus, specially, for any differentaible path $\{P_{\epsilon}:\epsilon\in\mathbb{R}\}\subsetneq\mathcal{M}$ that goes through $P_{0}$ at $\epsilon=0$, the parameter $\Psi(P;\pi)$ must be pathwise differentiable along this path at $P_{0}$ such that
$\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Psi(P_{\epsilon};\pi)\Big{|}_{\epsilon=0}=\mathrm{E}_{P_{0}}\Big{[}S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0}}(O)\dot{\ell}(O)\Big{]}$ (4)
where ˙$\dot{\ell}(O)$ is the score function uniquely determined by this parametric submodel introduced by this path.
However, when we turn to the optimal value functional $\Psi^{*}(P)=\Psi\big{(}P;\pi^{*}(P)\big{)}$. Intuitively, when there are no additional conditions, (4) may not be true anymore, as the optimal policy $\pi^{*}(P_{\epsilon})$ along this specific path may also contribute to this derivative when: (i) $\Psi\big{(}P;\pi\big{)}$ is sensitive to $\pi$; (ii) and $\pi^{*}(P)$ is sensitive to $P$. This means even when the EIF of $\Psi^{*}(P)$, $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}$, does exist, the chain rule of Gateaux differentials (Shapiro, 1990) implies
$\mathrm{E}_{P_{0}}\Big{[}S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}\big{|}_{P=P_{0}}(O)\dot{\ell}(O)\Big{]}$
$=$ $\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Psi\big{(}P_{\epsilon};\pi^{*}(P_{\epsilon})\big{)}\Big{|}_{\epsilon=0}$
$=$ $\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Psi\big{(}P_{\epsilon};\pi^{*}(P_{0})\big{)}\Big{|}_{\epsilon=0}+\mathrm{D}\Big{|}_{\frac{\mathrm{d}}{\mathrm{d}\epsilon}\pi^{*}(P_{\epsilon})|_{\epsilon=0}}\Psi\big{(}P_{0};\pi^{*}(P_{\epsilon})\big{)}\Big{|}_{\epsilon=0}$
$=$ $\mathrm{E}_{P_{0}}\Big{[}S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}(O)\dot{\ell}(O)\Big{]}+\mathds{D}\Big{|}_{\frac{\mathrm{d}}{\mathrm{d}\epsilon}\pi^{*}(P_{\epsilon})|_{\epsilon=0}}\Psi\big{(}P_{0};\pi^{*}(P_{\epsilon})\big{)}\Big{|}_{\epsilon=0}$
where $\mathds{D}|_{h}$ is the Gateaux differential along the specified direction $h$. Therefore, if the second term above is not zero, then
$S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}\big{|}_{P=P_{0}}\neq S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}.$
Our goal is to establish a neat expression for $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}=\psi(P)(\eta^{*},\zeta;O)$ with $\eta^{*}=\eta(\pi^{*})$ and $\zeta$ representing the nuisance functionals. Then we can easily construct an RAL estimator with this IF by solving $\widehat{\eta}^{*}$ in the following estimation equation
$\mathbb{P}_{N}\Big{[}\psi(P_{0})(\eta^{*},\zeta;O)-\mathrm{E}_{P(\eta^{*},\widehat{\zeta}(\eta^{*}))}[\psi(P_{0})(\eta^{*},\zeta;O)]\Big{]}=0,$
with
$\sqrt{N}\|\widehat{\zeta}(\eta^{*})-\zeta(P_{0})\|_{P_{0},\infty}={O}_{P_{0}}(1)$ (5)
From Chapter 3.3 of Tsiatis (2006), we know that $\widehat{\eta}^{*}$ from (5) is an RAL estimator with the IF $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}$ and thus achieves the semiparametric efficiency. Indeed, we will show that various well-known estimators in Section 4 all satisfy this display, thus are semiparametrically efficient.
3 Existence of Efficient Influence Function and Its Expression
To establish a sufficient condition for ensuring the EIF $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}$ exists, we present give the following additional regularity conditions on the distribution of $S$, $A$, and $R$ respectively. Let $\mathcal{S}$ and $\mathcal{A}$ be the supports of states and actions respectively. {assumption}[Regularity]
  • States: the state space $\mathcal{S}$ is compact and $S_{0}$ is not one-pointly distributed;
  • Actions: the action space $\mathcal{A}$ is finite and there exist positive constants $\underline{c}_{\pi}$ and $\overline{c}_{\pi}$ such that $\underline{c}_{\pi}\leq\inf_{\pi\in\mathcal{P}}\operatorname{essinf}_{(a,s)\in\mathcal{A}\times\mathcal{S}}\pi(a\mid s)\leq\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}$;
  • Rewards: $R$ is bounded;
All conditions in Assumption 3 is commonly-used in various MDP literature including OPE as well as Q-learning (for example, see the review papers Levine et al., 2020; Uehara et al., 2022; Shi, 2025), except the assumption that $0<\underline{c}_{\pi}\leq\inf_{\pi\in\mathcal{P}}\operatorname{essinf}_{(a,s)\in\mathcal{A}\times\mathcal{S}}\pi(a\mid s)\leq\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}<\infty$. However, this assumption is also quite mild: as one can see that if $\mathcal{S}$ is finite, then this assumption will be automatically satisfied, and such assumption ($\mathcal{S}$ is finite) is also quite commonly-used in OPE literature (see Xu et al., 2021; Shi et al., 2024b; Bian et al., 2024, for instance).
Our first result shows that when the optimal policy is unique and deterministic, then the EIF $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}$ exists and is equal to $S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}$ evaluated at $\pi=\pi^{*}$ exactly. We first state the unique and deterministic requirement formally in the following assumption.
{assumption}
[Unique Deterministic Optimal Policy] The optimal policy $\pi^{*}\in\mathcal{P}$ exists uniquely and it is deterministic such that
$\pi^{*}(P)(a\mid s)=\mathds{1}\Big{\{}a=\arg\max_{a^{\prime}\in\mathcal{A}}Q(P)(a^{\prime},s;\pi^{*})\Big{\}}\quad\text{ for }\quad s\in\mathcal{S}$
Indeed, in the context of infinite horizon MDPs, an optimal policy is often deterministic when the reward structure and transition dynamics are such that the value function’s maximization leads to a unique, deterministic action at each state (Puterman, 2014). Additionally, as discussed in Sutton et al. (1998), in many practical MDPs, especially those with deterministic or mildly stochastic transition dynamics and rewards, the optimal policy tends to be deterministic.
Theorem 3.1.
Suppose that Assumption 2.1, 2.1, 3, and 3 hold. Then the efficient influence function of $\Psi^{*}$ exists and satisfies
$S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}\big{|}_{P=P_{0}}=S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}.$
Our second result is that when the optimal policy is not unique such that there exists a large number of states $S$ that are indifferent (or incompatible) for at least two optimal polcies as stated in the following assumption, the influence function for $\Psi^{*}(P)$ does not exist at $P=P_{0}$.
{assumption}
[Unrestricted Optimal Rules] There exists at least one $\mathcal{P}\ni\pi^{\star}\neq\pi^{*}$ such that
$Q(P)\big{(}a,s;\pi^{\star}(P)\big{)}=Q(P)\big{(}a,s;\pi^{*}(P)\big{)}=\max_{\pi\in\mathcal{P}}Q(P)\big{(}a,s;\pi\big{)}$
and
$\mu\big{\{}s\in\mathcal{S}:\{a\in\mathcal{A}:\pi^{\star}(a\mid s)=\pi^{*}(a\mid s)\}\neq\varnothing\big{\}}>0.$
Assumption 3.1 is commonly referred to as Unrestricted Optimal Rules in the literature (see discussion in Robins & Rotnitzky, 2014), as it potentially leads to non-identification of the optimal policy. More significantly, this condition results in $\Psi^{*}(P_{0})$ no longer being pathwise differentiable, which has profound implications for OPE under optimal policies.
Theorem 3.2.
Suppose that Assumption 2.1, 2.1, 3, and 3.1 hold. Then $\Psi^{*}(P)$ does not have any influence function at $P=P_{0}$.
4 Examples
We present several commonly-used existing methods that achieve the semiparametric efficiency without parameterization or fixing the policies.
Example 4.1.
The first example is Shi et al. (2024b). Their Theorem 3 shows that conditional on the initial policy $\pi_{\text{old}}$ and given fixed policy $\pi$, the value difference estimator $\widehat{\eta}_{1}(\pi,\pi_{\text{old}})$ for
$\eta_{1}(\pi,\pi_{\text{old}}):=\sum_{a,s}\big{(}\pi(a\mid s)-\pi_{\text{old}}(a\mid s)\big{)}\big{(}Q(a,s;\pi_{\text{old}})-V(s;\pi_{\text{old}})\big{)}d(s;\pi_{\text{old}}),$
where $d(s^{\prime};\pi)=(1-\gamma)\sum_{a,s}\pi(a\mid s)\sum_{t\geq 0}\gamma^{t}f_{\sim\pi,t}(s^{\prime}\mid a,s)f_{+\infty}(s)$ is the integrated discounted visitation probability function, achieves semiparametric efficiency. Given the conditions in their Theorem 2, the learned policy, $\widehat{\pi}_{\text{new}}$, satisfies (5) as $O_{P_{0}}(\sqrt{N}(NT)^{-\frac{\kappa_{0}(2\alpha+1)}{\alpha+1}})+o_{P_{0}}(\sqrt{N}(NT)^{-1/2})=O_{P_{0}}(1)$. Therefore, under the conditions in Theorem 3.1, we establish that $\widehat{\eta}_{1}(\widehat{\pi}_{\text{new}},\pi_{\text{old}})$ still achieves the semiparametric efficiency conditional on $\pi_{\text{old}}$.
Example 4.2.
The second example is the widely-recognized Off-Policy Actor-Critic algorithm (see for example Degris et al., 2012; Wang et al., 2020; Xu et al., 2021), in which the policy $\pi$ is smooth enough and indexed by some parameter $w$. In this case, using the doubly robust off-policy actor-critic algorithm (Algorithm 1 in Xu et al., 2021) for the standard MDP, suppose we obtain the estimated policy as $\pi_{\widehat{w}}$ from Algorithm 1. If we regard their constructed doubly robust policy gradient error formula eq.(4) as a function of both $Q$ and $w$, denoted by $G_{\text{DR}}(Q;w)$. Then the estimation $\widehat{\eta}^{*}:=\mathbb{P}_{N}[\widehat{Q}(A_{0},S_{0};\pi_{\widehat{w}})\pi_{\widehat{w}}(A_{0}\mid S_{0})]$ with $\widehat{Q}$ derived from solving the estimation equation $\mathbb{P}_{NT}G_{\text{DR}}(Q;\widehat{w})=0$ will achieve the local semiparametric efficiency when the conditions in their Theorem 2 and our Theorem 3.1 are satisfied and the $\pi_{w}$ is correctly specified.
Example 4.3.
The last example is the combination of fitted-Q iteration (FQI) algorithm with doubly-robust FQE. When the deterministic optimal policy is unique, by plugging in the estimated policy learned by FQI into the doubly-robust FQE estimator, we obtain a semiparametrically efficient estimator. A simultaneous approach is also possible, which we leave for further study.
5 Final Remarks
In this paper, we have argued for the existence and characterization of the efficient influence function for $\eta(\pi^{*})$ under the canonical MDP framework widely recognized in statistical inference literature. However, our results can be straightforwardly extended to other OPE settings. For instance, both Theorem 3.1 and Theorem 3.2 can be directly applied to OPE scenarios when there exist mediators or instrumental variables (see Shi et al., 2024b; Xie et al., 2023; Liao et al., 2024).
We should note that we have not established the necessary and sufficient conditions for the existence of the EIF. We acknowledge that even when the optimal policy is not unique, the EIF may still exist, as demonstrated in various optimal treatment regime literatures (for instance Robins & Rotnitzky, 2014; Luedtke & Van Der Laan, 2016; Kosorok & Laber, 2019, although we do not find corresponding results for OPE). Such necessary and sufficient conditions remain an open question for future research.
Appendix A Technical Proofs
A.1 Preliminaries (under the same law $P$)
To prove the main results in Section 3, we first establish the connection between the distance between two policies and the value function via $Q$-functions or MIS functions. For notational clarity, we omit the underlying law $P$ for all functionals when they all follow the same probability distribution.
Lemma A.1 (Upper Bound).
Suppose that Assumption 2.1 and 2.1 hold. Then
$\big{\|}\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})\big{\|}_{1}\leq\frac{2\gamma}{1-\gamma}\mathrm{E}_{S\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S)\big{]}.$
Proof A.2.
Define the transition between $S_{t}\rightarrow S_{t+1}$ as
$K_{\pi}(s^{\prime}\mid s):=\int f(s^{\prime}\mid a,s)\pi(a\mid s)\,\mathrm{d}a.$
By the fact that $\{S_{t}\}_{t\geq 0}$ has been stationary under $b$, and we obtain
$\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})$ $=\frac{1-\gamma}{f_{0}(\cdot)}\sum_{t=0}^{+\infty}\gamma^{t}\big{(}f_{t}^{\pi_{2}}(\cdot)-f_{t}^{\pi_{1}}(\cdot)\big{)}=\frac{1-\gamma}{f(\cdot)}\sum_{t=0}^{+\infty}\big{(}\gamma^{t}K_{\pi_{2}}^{t}-\gamma^{t}K_{\pi_{1}}^{t}\big{)}f_{0}(\cdot)$
$=\frac{1-\gamma}{f_{0}(\cdot)}\gamma(I-\gamma K_{\pi_{2}})^{-1}(K_{\pi_{2}}-K_{\pi_{1}})(I-\gamma K_{\pi_{1}})^{-1}f_{0}(\cdot)$
$=\gamma(I-\gamma K_{\pi_{2}})^{-1}(K_{\pi_{2}}-K_{\pi_{1}})\frac{(1-\gamma)(I-\gamma K_{\pi_{1}})^{-1}f_{0}(\cdot)}{f_{0}(\cdot)}$
$=\gamma(I-K_{\pi_{2}})^{-1}(K_{\pi_{2}}-K_{\pi_{1}})\omega(\cdot;\pi_{1}).$
Note that $K_{\pi}$ is a probability kernel, we have
$\|(I-\gamma K_{\pi_{2}})^{-1}\|_{1}\leq\sum_{t\geq 0}\gamma^{t}\|K_{\pi_{2}}\|_{1}^{t}\leq\sum_{t\geq 0}\gamma^{t}1=(1-\gamma)^{-1}$
which implies
$\big{\|}\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})\big{\|}_{1}$ $\leq\gamma\big{\|}(I-K_{\pi_{2}})^{-1}\big{\|}_{1}\big{\|}(K_{\pi_{2}}-K_{\pi_{1}})\omega(\cdot;\pi_{1})\big{\|}_{1}$
$\leq\frac{\gamma}{1-\gamma}\big{\|}(K_{\pi_{2}}-K_{\pi_{1}})\omega(\cdot;\pi_{1})\big{\|}_{1}$
$=\frac{\gamma}{1-\gamma}\int\mathrm{d}s^{\prime}\Big{|}(K_{\pi_{2}}-K_{\pi_{1}})(s^{\prime}\mid s)\omega(s;\pi_{1})\mathrm{d}s\Big{|}$
$\leq\frac{\gamma}{1-\gamma}\int f(s^{\prime}\mid a,s)\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\omega(s;\pi_{1})\,\mathrm{d}(s^{\prime},a,s)$
$=\frac{\gamma}{1-\gamma}\int\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\omega(s;\pi_{1})\,\mathrm{d}(a,s)=\frac{2\gamma}{1-\gamma}\mathrm{E}_{S\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S)\big{]},$
by the definition of the total variation distance $\operatorname{TV}(\pi_{2}\|\pi_{1})$.
Lemma A.3 (Lower Bound).
Suppose the conditions in Lemma A.1 hold. If
$\underline{c}_{\pi}\leq\operatorname{essinf}_{\pi\in\mathcal{P}}\operatorname{essinf}_{(a,s)\in\mathcal{A}\times\mathcal{S}}\pi(a\mid s)\leq\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}$
then
$\big{|}\omega(s;\pi_{2})-\omega(s;\pi_{1})\big{|}\geq\frac{2\gamma\sqrt{\overline{c}_{\pi}^{-1}\underline{c}_{\pi}\mathrm{E}_{S^{\prime}\sim\omega(s;\pi_{1})}\big{[}\chi^{2}(\pi_{2}\|\pi_{1})(S^{\prime})\big{]}}}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}\sqrt{f(s)}$
for any $s\in\mathcal{S}$.
Proof A.4.
Similar as in the proof of Lemma A.1, we have the identity
$\omega(s;\pi_{2})-\omega(s;\pi_{1})=\gamma(I-\gamma K_{\pi_{2}})^{-1}(K_{\pi_{2}}-K_{\pi_{1}})\omega(s;\pi_{1}).$
For any positive density $h$,
$\big{(}(I-\gamma K_{\pi_{2}})^{-1}h\big{)}(s)=\sum_{t=0}^{\infty}\gamma^{t}K_{\pi_{2}}^{t}h(s)=\bigg{(}h+\sum_{t=1}^{\infty}\gamma^{t}K_{\pi_{2}}^{t}h\bigg{)}(s)\geq 1\cdot h(s)=h(s)$
holds pointwisely for any $s$. Thus, an initial lower bound can be obtained as
$\omega(s;\pi_{2})-\omega(s;\pi_{1})$ $=\gamma(I-\gamma K_{\pi_{2}})^{-1}(K_{\pi_{2}}-K_{\pi_{1}})\omega(s;\pi_{1})$
$\geq\gamma(K_{\pi_{2}}-K_{\pi_{1}})\omega(s;\pi_{1})$
$=\gamma\int f(s\mid a,s_{\star})\big{[}\pi_{2}(a\mid s_{\star})-\pi_{1}(a\mid s_{\star})\big{]}\omega(s_{\star};\pi_{1})\,\mathrm{d}a\,\mathrm{d}s_{\star}.$
To include the Kullback-Leibler divergence in the lower bound, we first apply reverse Pinsker inequalities (for example, see Cattiaux & Guillin, 2009) as
$\operatorname{KL}(\pi_{2}\|\pi_{1})(s)\leq\frac{2}{\underline{c}_{s}}\operatorname{TV}^{2}(\pi_{2}\|\pi_{1})(s)=\frac{1}{2\underline{c}_{s}}\bigg{(}\int\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\,\mathrm{d}a\bigg{)}^{2}.$ (6)
The second tool for obtaining the lower bound is using the reverse Cauchy inequality (see the result for integrals in Corollary 6.1 of Aldaz et al., 2015). The bounds for the ratio are
$\frac{\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}{\sqrt{\overline{c}_{\pi}}}$ $\leq\frac{\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}{\sqrt{\overline{c}_{\pi}}}$
$\leq\frac{|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}{\sqrt{\pi_{1}(a\mid s)}}\leq\frac{\overline{c}_{\pi}}{\sqrt{\underline{c}_{\pi}}},$
and thus we obtain
$\bigg{(}\int f(s\mid a,s_{\star})\big{[}\pi_{2}(a\mid s_{\star})-\pi_{1}(a\mid s_{\star})\big{]}\omega(s_{\star};\pi_{1})\,\mathrm{d}a\,\mathrm{d}s_{\star}\bigg{)}^{2}$
$=$ $\bigg{(}\int\omega(s_{\star};\pi_{1})f(s\mid a,s_{\star})\big{[}\pi_{2}(a\mid s_{\star})-\pi_{1}(a\mid s_{\star})\big{]}\,\mathrm{d}a\,\mathrm{d}s_{\star}\bigg{)}^{2}$
$\geq$ $\frac{4\underline{c}_{\pi}^{-1/2}\overline{c}_{\pi}\times\overline{c}_{\pi}^{-1/2}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}{\big{(}\underline{c}_{\pi}^{-1/2}\overline{c}_{\pi}+\overline{c}_{\pi}^{-1/2}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|\big{)}^{2}}$
$\int\omega(s_{\star};\pi_{1})f(s\mid a,s_{\star})\bigg{(}\frac{\pi_{2}(a\mid s_{\star})-\pi_{1}(a\mid s_{\star})}{\sqrt{\pi_{1}(a\mid s_{\star})}}\bigg{)}^{2}\,\mathrm{d}a\,\mathrm{d}s_{\star}\int\omega(s_{\star};\pi_{1})f(s\mid a,s_{\star})\times\big{[}\sqrt{\pi_{1}(a\mid s_{\star})}\big{]}^{2}\,\mathrm{d}a\,\mathrm{d}s_{\star}$
$=$ $\frac{4\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}{\underline{c}_{\pi}^{-1/2}\overline{c}_{\pi}^{1/2}\big{(}\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|\big{)}^{2}}\times f(s)$
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\int\omega(s_{\star};\pi_{1})f(s\mid a,s_{\star})\bigg{(}\frac{\pi_{2}(a\mid s_{\star})-\pi_{1}(a\mid s_{\star})}{\sqrt{\pi_{1}(a\mid s_{\star})}}\bigg{)}^{2}\,\mathrm{d}a\,\mathrm{d}s_{\star}.$
By using the definition of chi-square divergence, the integral in the above expression can be rewritten as
$\int\omega(s_{\star};\pi_{1})f(s\mid a,s_{\star})\bigg{(}\frac{\pi_{2}(a\mid s_{\star})-\pi_{1}(a\mid s_{\star})}{\sqrt{\pi_{1}(a\mid s_{\star})}}\bigg{)}^{2}\,\mathrm{d}a\,\mathrm{d}s_{\star}$
$=$ $\int\omega(s_{\star};\pi_{1})f(s\mid a,s_{\star})\pi_{1}(a\mid s_{\star})\bigg{(}\frac{\pi_{2}(a\mid s_{\star})}{\pi_{1}(a\mid s_{\star})}-1\bigg{)}^{2}\,\mathrm{d}a\,\mathrm{d}s_{\star}=\mathrm{E}_{S^{\prime}\sim\omega(s;\pi_{1})}\big{[}\chi^{2}(\pi_{2}\|\pi_{1})(S^{\prime})\big{]},$
which leads to the result in the lemma.
To our best knowledge, our novel result in Lemma A.3 is the first one to find the lower bound for the divergence between different policies, while there exists study on the upper bounds for $\omega(s;\pi_{2})-\omega(s;\pi_{1})$ (for example, see Achiam et al., 2017; Huang & Jiang, 2024; Krishnamurthy et al., 2025). We does not compare the tightness of these upper bounds with ours here, as ours is enough for obtaining our results. Additionally, the lower bound in Lemma A.3, which is first studied to our best knowledge, is more important than the these upper bounds, as it can help connect the difference between policies with the evalution $Q$-function. To do this, we introduce some helpful tools and define the advantage function $\mathbb{A}(a,s;\pi)$ as
$\mathbb{A}(a,s;\pi):=Q(a,s;\pi)-V(s;\pi).$
The advantage function $\mathbb{A}(a,s;\pi)$ will help us connect the difference $|\omega(s;\pi_{2})-\omega(s;\pi_{1})|$ with $|Q(a,s;\pi_{2})-Q(a,s;\pi_{1})|$ if both $\pi_{1}$ and $\pi_{2}$ come from greedy policy (may not be unique).
Lemma A.5.
Suppose the conditions in Lemma A.3 hold. If $S_{0}$ is not one-pointly distributed and the action space $\mathcal{A}$ is finite, then the total variation distance between $\pi_{2}$ and $\pi_{1}$ can be both upper and lower bounded by the difference of their corresponding $Q$-functions via (10).
Proof A.6.
We expand the identity in Lemma A.7 as
$\eta(\pi_{2})-\eta(\pi_{1})$ $=\frac{1}{1-\gamma}\mathrm{E}_{S\sim\omega(S_{0};\pi_{2})}\big{[}\mathrm{E}_{A\sim\pi_{2}(\cdot\mid S)}\mathbb{A}(A,S;\pi_{1})\big{]}$
$=\frac{1}{1-\gamma}\int\omega(s;\pi_{2})\,\mathrm{d}s\int\pi_{2}(a\mid s)\mathbb{A}(a,s;\pi_{1})\,\mathrm{d}a$
$=\frac{1}{1-\gamma}\int\omega(s;\pi_{2})\,\mathrm{d}s\int\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}Q(a,s;\pi_{1})\,\mathrm{d}a$
by noting that $S_{0}$ is already stationary. Similarly,
$\eta(\pi_{1})-\eta(\pi_{2})=\frac{1}{1-\gamma}\int\omega(s;\pi_{1})\,\mathrm{d}s\int\big{[}\pi_{1}(a\mid s)-\pi_{2}(a\mid s)\big{]}Q(a,s;\pi_{2})\,\mathrm{d}a,$
which implies
$2\big{[}\eta(\pi_{2})-\eta(\pi_{1})\big{]}=\frac{1}{1-\gamma}\int\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\Big{(}\omega(s;\pi_{2})Q(a,s;\pi_{1})-\omega(s;\pi_{1})Q(a,s;\pi_{2})\Big{)}\,\mathrm{d}(a,s).$
Therefore, we obtain the difference between value functions under different policies as
$\eta(\pi_{2})-\eta(\pi_{1})$ $=\frac{1}{2(1-\gamma)}\int\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\omega(s;\pi_{2})\Big{(}Q(a,s;\pi_{1})-Q(a,s;\pi_{2})\Big{)}\,\mathrm{d}(a,s)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{2(1-\gamma)}\int\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\Big{(}\omega(s;\pi_{2})-\omega(s;\pi_{1})\Big{)}Q(a,s;\pi_{2})\,\mathrm{d}(a,s).$
The above expression together with the following decomposition
$\eta(\pi_{2})-\eta(\pi_{1})$
$=$ $\mathrm{E}\big{[}V(S_{0};\pi_{2})-V(S_{0};\pi_{1})\big{]}$
$=$ $\int f(s)\,\mathrm{d}s\int\Big{(}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\Big{)}\pi_{2}(a\mid s)\,\mathrm{d}a+\int f(s)\,\mathrm{d}s\int Q(a,s;\pi_{1})\Big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\Big{)}\,\mathrm{d}a,$
we can rewrite the difference as
$\frac{1}{2(1-\gamma)}\int\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\Big{(}\omega(s;\pi_{2})-\omega(s;\pi_{1})\Big{)}Q(a,s;\pi_{2})\,\mathrm{d}(a,s)$ (7)
$=$ $\int f(s)Q(a,s;\pi_{1})\Big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\Big{)}\,\mathrm{d}(a,s)$
$+\int\bigg{(}f(s)\pi_{2}(a\mid s)+\frac{1}{2(1-\gamma)}\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\omega(s;\pi_{2})\bigg{)}\Big{(}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\Big{)}\mathrm{d}(a,s).$
By using the fact that the action space $\mathcal{A}$ is finite, the left hand in (7) can be lower bounded by
$\bigg{|}\frac{1}{2(1-\gamma)}\int\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\Big{(}\omega(s;\pi_{2})-\omega(s;\pi_{1})\Big{)}Q(a,s;\pi_{2})\,\mathrm{d}(a,s)\bigg{|}$
$\geq$ $\frac{1}{2(1-\gamma)}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\bigg{|}\int\Big{(}\omega(s;\pi_{2})-\omega(s;\pi_{1})\Big{)}Q(a,s;\pi_{2})\,\mathrm{d}(a,s)\bigg{|}$
$\geq$ $\frac{|\mathcal{A}|\underline{c}_{R}}{2(1-\gamma)}\bigg{|}\int\Big{(}\omega(s;\pi_{2})-\omega(s;\pi_{1})\Big{)}\,\mathrm{d}s\bigg{|}=\frac{|\mathcal{A}|\underline{c}_{R}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}}{2(1-\gamma)}\big{\|}\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})\big{\|}_{1}.$
Similarly, for finding a suitable upper bound for second term of the right hand in (7), we use the fact $\|\pi_{2}\|_{\infty}\leq\overline{c}_{\pi}$ and obtain
$\bigg{|}\int\bigg{(}f(s)\pi_{2}(a\mid s)+\frac{1}{2(1-\gamma)}\big{[}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{]}\omega(s;\pi_{2})\bigg{)}\Big{(}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\Big{)}\mathrm{d}(a,s)\bigg{|}$
$\leq$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}\int\bigg{(}f(s)\pi_{2}(a\mid s)+\frac{1}{2(1-\gamma)}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\omega(s;\pi_{2})\bigg{)}\mathrm{d}(a,s)$
$=$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}\bigg{(}1+\frac{1}{2(1-\gamma)}\int\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\omega(s;\pi_{2})\,\mathrm{d}(a,s)\bigg{)}$
$=$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}\bigg{(}1+\frac{1}{(1-\gamma)}\mathrm{E}_{S\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S)\big{]}\bigg{)}.$
Plugging the above three inequalities back into (7), we obtain
$\frac{|\mathcal{A}|\underline{c}_{R}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}}{2(1-\gamma)}\big{\|}\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})\big{\|}_{1}$ (8)
$\leq$ $\int f(s)Q(a,s;\pi_{1})\Big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\Big{)}\,\mathrm{d}(a,s)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}\bigg{(}1+\frac{1}{(1-\gamma)}\mathrm{E}_{S\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S)\big{]}\bigg{)}.$
To avoid the square root in Lemma A.3, we apply the Pinsker inequalities (for example, Proposition 2.2.9 in Duchi, 2015) as
${\chi}^{2}(\pi_{2}\|\pi_{1})\geq\operatorname{KL}(\pi_{2}\|\pi_{1})\geq 2\operatorname{TV}^{2}(\pi_{2}\|\pi_{1}),$
which implies the following pointwise inequality
$\big{|}\omega(s;\pi_{2})-\omega(s;\pi_{1})\big{|}$ $\geq\frac{2\gamma\sqrt{\overline{c}_{\pi}^{-1}\underline{c}_{\pi}\mathrm{E}_{S^{\prime}\sim\omega(s;\pi_{1})}\big{[}\chi^{2}(\pi_{2}\|\pi_{1})(S^{\prime})\big{]}}}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}\sqrt{f(s)}$
$\geq\frac{2\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}f(s)}}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}\mathrm{E}_{S^{\prime}\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S^{\prime})\big{]}.$
and furthermore
$\big{\|}\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})\big{\|}_{1}\geq\frac{2\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}\mathrm{E}_{S^{\prime}\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S^{\prime})\big{]},$ (9)
where $\operatorname{BC}(S_{0},U(\mathcal{S}))\in[0,1]$ is the Bhattacharyya coefficient between the stationary distribution of $S$ and the uniform distribution on $\mathcal{S}$ and it is strictly positive by $S_{0}$ is not a one-point distribution (see discussion in Ali & Silvey, 1966).
It remains to upper bound the first term of the right hand in (7). We rewrite the first term of the right hand in (7) as
$\int f(s)Q(a,s;\pi_{1})\Big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\Big{)}\,\mathrm{d}(a,s)$
$=$ $\int f(s)\,\mathrm{d}s\bigg{(}\int Q(a,s;\pi_{1})\pi_{2}(a\mid s)\,\mathrm{d}a-V(s;\pi_{2})+V(s;\pi_{2})-V(s;\pi_{1})\bigg{)}$
$=$ $\int f(s)\,\mathrm{d}s\bigg{(}\int\big{[}Q(a,s;\pi_{1})-Q(a,s;\pi_{2})\big{]}\pi_{2}(a\mid s)\,\mathrm{d}a+V(s;\pi_{2})-V(s;\pi_{1})\bigg{)}.$
Thus,
$\bigg{|}\int f(s)Q(a,s;\pi_{1})\Big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\Big{)}\,\mathrm{d}(a,s)\bigg{|}$
$\leq$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}\int f(s)\,\mathrm{d}s\,\pi_{2}(a\mid s)\,\mathrm{d}a+\big{\|}V(s;\pi_{2})-V(s;\pi_{1})\big{\|}_{\infty}\int f(s)\,\mathrm{d}s$
$=$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}+\big{\|}V(s;\pi_{2})-V(s;\pi_{1})\big{\|}_{\infty}.$
For $\|V(s;\pi_{2})-V(s;\pi_{1})\|_{\infty}$, introduce the Bellman operator $\mathbb{T}^{\pi}$ such that
$(\mathbb{T}^{\pi}V)(s;\pi)=\int\pi(a\mid s)\,\mathrm{d}a\bigg{(}\mathrm{E}[R\mid A=a,S=s]+\gamma\int V(s^{\prime};\pi)f(s^{\prime}\mid a,s)\,\mathrm{d}s^{\prime}\bigg{)}.$
then we have identity $V(s;\pi)=\mathbb{T}^{\pi}V(s;\pi)$ and it is a contraction operator with coefficient $\gamma$ (see Shi, 2025, for example). Therefore,
$\big{\|}V(s;\pi_{2})-V(s;\pi_{1})\big{\|}_{\infty}$ $=\big{\|}\mathbb{T}^{\pi_{2}}V(s;\pi_{2})-\mathbb{T}^{\pi_{2}}V(s;\pi_{1})+\mathbb{T}^{\pi_{2}}V(s;\pi_{1})-\mathbb{T}^{\pi_{1}}V(s;\pi_{1})\big{\|}_{\infty}$
$\leq\gamma\big{\|}V(s;\pi_{2})-V(s;\pi_{1})\big{\|}_{\infty}+\big{\|}\mathbb{T}^{\pi_{2}}V(s;\pi_{1})-\mathbb{T}^{\pi_{1}}V(s;\pi_{1})\big{\|}_{\infty}$
$=\gamma\big{\|}V(s;\pi_{2})-V(s;\pi_{1})\big{\|}_{\infty}+\big{\|}\mathbb{T}^{\pi_{2}}V(s;\pi_{1})-\mathbb{T}^{\pi_{1}}V(s;\pi_{1})\big{\|}_{\infty}$
with the second term above is bounded by
$\big{\|}\mathbb{T}^{\pi_{2}}V(s;\pi_{1})-\mathbb{T}^{\pi_{1}}V(s;\pi_{1})\big{\|}_{\infty}$ $=\sup_{s}\bigg{|}\int\big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{)}Q(a,s;\pi_{1})\,\mathrm{d}a\bigg{|}$
$\leq\sup_{s\in\mathcal{S}}\sup_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\|Q(a,s;\pi_{1})\|_{\infty}$
$\leq\sup_{s\in\mathcal{S}}\sup_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\frac{\overline{c}_{R}}{1-\gamma}.$
This leads to the bounds for $\|V(s;\pi_{2})-V(s;\pi_{1})\|_{\infty}$ and the first term of the right hand in (7) as
$\big{\|}V(s;\pi_{2})-V(s;\pi_{1})\big{\|}_{\infty}\leq\sup_{s\in\mathcal{S}}\sup_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}\frac{\overline{c}_{R}}{(1-\gamma)^{2}}$
and
$\bigg{|}\int f(s)Q(a,s;\pi_{1})\Big{(}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\Big{)}\,\mathrm{d}(a,s)\bigg{|}$
$\leq$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}+\frac{\overline{c}_{R}\sup_{s\in\mathcal{S}}\sup_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}}{(1-\gamma)^{2}}$
Applying this bound with (9) on the inequality (8), we can give a upper bound for $\mathrm{E}_{S\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S)\big{]}$ as
$\frac{|\mathcal{A}|\underline{c}_{R}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}}{2(1-\gamma)}$ (10)
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{2\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}|\pi_{2}(a\mid s)-\pi_{1}(a\mid s)|}\mathrm{E}_{S^{\prime}\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S^{\prime})\big{]}$
$\leq$ $\frac{|\mathcal{A}|\underline{c}_{R}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}}{2(1-\gamma)}\big{\|}\omega(\cdot;\pi_{2})-\omega(\cdot;\pi_{1})\big{\|}_{1}$
$\leq$ $\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}+\frac{\overline{c}_{R}\sup_{s\in\mathcal{S}}\sup_{a\in\mathcal{A}}\big{|}\pi_{2}(a\mid s)-\pi_{1}(a\mid s)\big{|}}{(1-\gamma)^{2}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\big{\|}Q(a,s;\pi_{2})-Q(a,s;\pi_{1})\big{\|}_{\infty}\bigg{(}1+\frac{1}{(1-\gamma)}\mathrm{E}_{S\sim\omega(s;\pi_{1})}\big{[}\operatorname{TV}(\pi_{2}\|\pi_{1})(S)\big{]}\bigg{)},$
which completes our proof.
Now we can obtain the lower bound and upper bound for the distance of policies with respect to the $Q$-functions.
A.2 Decomposing the Gateaux differential of $\Psi^{*}(P_{\epsilon})$
For any $P\in\mathcal{M}$ and fixed policy $\pi\in\mathcal{{P}}$,
$\Psi^{*}(P)-\Psi(P;\pi)$
$=$ $\Psi^{*}(P)-\mathrm{E}_{P}\big{[}Q(P)(A,S;\pi)\pi(A\mid S)\big{]}$
$=$ $\mathrm{E}_{P}\big{[}Q(P)(A,S;\pi^{*}(P))\pi^{*}(P)(A\mid S)\big{]}-\mathrm{E}_{P}\big{[}Q(P)(A,S;\pi)\pi(A\mid S)\big{]}$
$=$ $\mathrm{E}_{P}\big{[}Q(P)\big{(}A,S;\pi^{*}(P)\big{)}\big{(}\pi^{*}(P)(A\mid S)-\pi(A\mid S)\big{)}\big{]}+\mathrm{E}_{P}\Big{[}\Big{(}Q(P)\big{(}A,S;\pi^{*}(P)\big{)}-Q(P)(A,S;\pi)\Big{)}\pi(A\mid S)\Big{]}$
$=$ $\mathrm{E}_{P}\big{[}Q(P)\big{(}A,S;\pi^{*}(P)\big{)}\big{(}\pi^{*}(P)(A\mid S)-\pi(A\mid S)\big{)}\big{]}+\mathrm{E}_{P}\big{[}\Delta(P)(\pi;A,S)\pi(A\mid S)\big{]}$
which implies
$\Psi^{*}(P_{\epsilon})-\Psi^{*}(P_{0})$
$=$ $\mathrm{E}_{P_{\epsilon}}\big{[}Q(P_{\epsilon})\big{(}A,S;\pi^{*}(P_{\epsilon})\big{)}\big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi(A\mid S)\big{)}\big{]}-\mathrm{E}_{P_{0}}\big{[}Q(P_{0})\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}\big{]}$
$~{}~{}~{}~{}~{}~{}+\mathrm{E}_{P_{\epsilon}}\big{[}\Delta(P_{\epsilon})(\pi;A,S)\pi(A\mid S)\big{]}-\mathrm{E}_{P_{0}}\big{[}\Delta(P_{0})(\pi;A,S)\pi(A\mid S)\big{]}+\Psi(P_{\epsilon};\pi)-\Psi(P_{0};\pi).$
It is well known that $\Psi(P_{\epsilon};\pi)-\Psi(P_{0};\pi)$ is pathwise differentiable for fixed $\pi$ with the exactly same derivates shown in (4). In addition,
$\mathrm{E}_{P_{\epsilon}}\big{[}Q(P_{\epsilon})\big{(}A,S;\pi^{*}(P_{\epsilon})\big{)}\big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi(A\mid S)\big{)}\big{]}-\mathrm{E}_{P_{0}}\big{[}Q(P_{0})\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}\big{]}$
$=$ $\mathrm{E}_{P_{\epsilon}}\big{[}\Delta(P_{\epsilon})(\pi;A,S)\big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi(A\mid S)\big{)}\big{]}-\mathrm{E}_{P_{\epsilon}}\big{[}\Delta(P_{0})(\pi;A,S)\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}\big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathrm{E}_{P_{\epsilon}}\big{[}Q(P_{0})(\pi;A,S)\big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)\big{)}\big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\big{(}\mathrm{E}_{P_{\epsilon}}-\mathrm{E}_{P_{0}}\big{)}\big{[}Q(P_{0})\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}\big{]}.$
and
$\mathrm{E}_{P_{\epsilon}}\big{[}\Delta(P_{\epsilon})(\pi;A,S)\pi(A\mid S)\big{]}-\mathrm{E}_{P_{0}}\big{[}\Delta(P_{0})(\pi;A,S)\pi(A\mid S)\big{]}$
$=$ $\mathrm{E}_{P_{\epsilon}}\big{[}\big{(}\Delta(P_{\epsilon})(\pi;A,S)-\Delta(P_{0})(\pi;A,S)\big{)}\pi(A\mid S)\big{]}+\big{(}\mathrm{E}_{P_{\epsilon}}-\mathrm{E}_{P_{0}}\big{)}\big{[}\Delta(P_{0})(\pi;A,S)\pi(A\mid S)\big{]}.$
Since $\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}$ and $\|Q\|_{\infty},\|\Delta\|_{\infty}\leq(1-\gamma)^{-1}\overline{c}_{R}$, both
$\big{(}\mathrm{E}_{P_{\epsilon}}-\mathrm{E}_{P_{0}}\big{)}\big{[}Q(P_{0})\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}\big{]}$
and
$\big{(}\mathrm{E}_{P_{\epsilon}}-\mathrm{E}_{P_{0}}\big{)}\big{[}\Delta(P_{0})(\pi;A,S)\pi(A\mid S)\big{]}$
are pathwise differentiable for any differentiable path $P_{\epsilon}$ such that (for example, see Theorem 25.81 in Van Der Vaart, 2000)
$\big{(}\mathrm{E}_{P_{\epsilon}}-\mathrm{E}_{P_{0}}\big{)}\big{[}Q(P_{0})\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}+\Delta(P_{0})(\pi;A,S)\pi(A\mid S)\big{]}=o_{P_{0}}(\epsilon).$
Therefore, we have the following decomposition
$\Psi^{*}(P_{\epsilon})-\Psi^{*}(P_{0})$ (11)
$=$ $\underbrace{\mathrm{E}_{P_{\epsilon}}\big{[}\Delta(P_{\epsilon})(\pi;A,S)\big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi(A\mid S)\big{)}\big{]}-\Delta(P_{0})(\pi;A,S)\big{(}\pi^{*}(P_{0})(A\mid S)-\pi(A\mid S)\big{)}\big{]}}_{=:\mathbbl{\Delta}_{1}(\epsilon;\pi)}$
$+\underbrace{\mathrm{E}_{P_{\epsilon}}\big{[}Q(P_{0})(\pi;A,S)\big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)\big{)}\big{]}}_{=:\mathbbl{\Delta}_{2}(\epsilon;\pi)}$
$+\underbrace{\mathrm{E}_{P_{\epsilon}}\big{[}\big{(}\Delta(P_{\epsilon})(\pi;A,S)-\Delta(P_{0})(\pi;A,S)\big{)}\pi(A\mid S)\big{]}}_{=:\mathbbl{\Delta}_{3}(\epsilon;\pi)}+\Psi(P_{\epsilon};\pi)-\Psi(P_{0};\pi)+o_{P_{0}}(\epsilon).$
The decomposition (11) implies the key to analyze the Gateaux differential lies on two crucial differences:
  • $\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)$ shown in $\mathbbl{\Delta}_{1}(\epsilon;\pi)$ and $\mathbbl{\Delta}_{2}(\epsilon;\pi)$;
  • $\Delta(P_{\epsilon})(\pi;A,S)-\Delta(P_{0})(\pi;A,S)$ shown in $\mathbbl{\Delta}_{2}(\epsilon;\pi)$ and $\mathbbl{\Delta}_{3}(\epsilon;\pi)$.
Notice that
$\big{\|}Q(P_{\epsilon})\big{(}a,s;\pi^{*}(P_{\epsilon})\big{)}-Q(P_{0})\big{(}a,s;\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},\infty}$ (12)
$=$ $\Big{\|}\sup_{\pi\in\mathcal{P}}Q(P_{\epsilon})(a,s;\pi)-\sup_{\pi\in\mathcal{P}}Q(P_{0})(a,s;\pi)\Big{\|}_{P_{0},\infty}$
$\leq$ $\sup_{\pi\in\mathcal{P}}\big{\|}Q(P_{\epsilon})(a,s;\pi)-Q(P_{0})(a,s;\pi)\big{\|}_{P_{0},\infty}$
$=$ $\sup_{\pi\in\mathcal{P}}\bigg{\|}\Big{(}\mathrm{E}_{P_{\epsilon}}^{\sim\pi}-\mathrm{E}_{P_{0}}^{\sim\pi}\Big{)}\bigg{[}\sum_{k=0}^{+\infty}\gamma^{k}R_{t+k}\mid A_{t}=a,S_{t}=s\bigg{]}\bigg{\|}_{P_{0},\infty}$
$\leq$ $\sup_{\pi\in\mathcal{P}}\bigg{\|}\Big{(}\mathrm{E}_{P_{\epsilon}}^{\sim\pi}-\mathrm{E}_{P_{0}}^{\sim\pi}\Big{)}\bigg{[}\sum_{k=0}^{+\infty}\gamma^{k}\overline{c}_{R}\mid A_{t}=a,S_{t}=s\bigg{]}\bigg{\|}_{P_{0},\infty}$
$\leq$ $\frac{\overline{c}_{R}}{1-\gamma}\sup_{\pi\in\mathcal{P}}\Big{\|}\mathrm{E}_{P_{\epsilon}}^{\sim\pi}-\mathrm{E}_{P_{0}}^{\sim\pi}\Big{\|}_{P_{0},\infty}=o_{P_{0}}(\epsilon),$
is well-behaved for any differentiable path $P_{\epsilon}$, thus the same for $\Delta(P_{\epsilon})(\pi;A,S)-\Delta(P_{0})(\pi;A,S)=Q(P_{\epsilon})\big{(}a,s;\pi^{*}(P_{\epsilon})\big{)}-Q(P_{0})\big{(}a,s;\pi^{*}(P_{0})\big{)}$. All of remaining task is bounding $\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)$, we will finish this in the next subsection.
A.3 Bound on $\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)$
Let us talk with $\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)$ first. By the definition and Assumption 3, we have
$\pi^{*}(P_{\epsilon})(a\mid s)=\arg\max_{\pi\in\mathcal{P}}Q(P_{\epsilon})(a,s;\pi)\quad\text{ and }\quad\pi^{*}(P_{0})(a\mid s)=\arg\max_{\pi\in\mathcal{P}}Q(P_{0})(a,s;\pi).$
We need to carefully apply the result in Lemma A.5 as they come from different underlying distributions. Dor the expectation of the total variation distance between $\pi^{*}(P_{\epsilon})$ and $\pi^{*}(P_{0})$, considering the following regular submodel (satisfying the differentiation property in quadratic mean, see Section 2.53 in Van Der Vaart, 2000) such that
$\mathrm{d}P_{S,\epsilon}-\mathrm{d}P_{S,0}=\big{(}1+\epsilon h_{S}(S)\big{)}\mathrm{d}P_{S,0},\quad\text{ where }\quad\mathrm{E}_{P_{0}}[h_{S}(S)]=0;$ (13)
and $\mathrm{d}P_{R,\epsilon}-\mathrm{d}P_{R,0}=\big{(}1+\epsilon h_{R}(R\mid A,S)\big{)}\mathrm{d}P_{R,0},\quad\text{ where }\quad\mathrm{E}_{P_{0}}[h_{R}(R\mid A,S)]=0.$
Now assume there exist $(a,\mathcal{S}_{\star})\in\mathcal{A}\times 2^{\mathcal{S}}$ and $\delta\in(0,\overline{c}_{\pi}]$ such that $\mu(\mathcal{S}\,\setminus\,\mathcal{S}_{\star})=0$ and
$\inf_{s\in\mathcal{S}_{\star}}\big{|}\pi^{*}(P_{\epsilon})(a\mid s)-\pi^{*}(P_{0})(a\mid s)\big{|}:=\delta,$ (14)
then the expectation of the total variation distance between $\pi^{*}(P_{\epsilon})$ and $\pi^{*}(P_{0})$ in (10) is lower bounded by
$\frac{1}{2}\int\big{|}\pi^{*}(P_{\epsilon})(a\mid s)-\pi^{*}(P_{0})(a\mid s)\big{|}\,\mathrm{d}a\,\omega(P_{0})\big{(}s;\pi^{*}(P_{0})\big{)}\,\mathrm{d}s\geq\delta/2,$
where the assumption that $\mathcal{A}$ is finite is necessary hre. Therefore, the lower bound in (10) can be refined as
$\frac{|\mathcal{A}|\underline{c}_{R}\delta}{2(1-\gamma)}\big{\|}\omega(P_{\epsilon})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},1}\geq\frac{|\mathcal{A}|\underline{c}_{R}\delta}{2(1-\gamma)}\frac{2\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\delta}\frac{\delta}{2},$
i.e.,
$\big{\|}\omega(P_{\epsilon})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},1}\geq\frac{\delta\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{1+\overline{c}_{\pi}^{1/2}}.$ (15)
Now, we will show that the above expression will lead to a contradiction. We use the decomposition in Lemma A.8 and switch the order $\pi^{*}(P_{\epsilon})$ and $\pi^{*}(P_{0})$, this lead to the following two identities
[…]
$=$ $\mathrm{E}_{S\sim\omega(P_{\epsilon})(S;\pi^{*}(P_{\epsilon})),A\sim\pi^{*}(P_{\epsilon})}\delta_{f}(S^{\prime},A,S)-\mathrm{E}_{S\sim\omega(P_{0})(S;\pi^{*}(P_{0})),A\sim\pi^{*}(P_{0})}\bigg{[}\frac{\pi^{*}(P_{\epsilon})(A\mid S)}{\pi^{*}(P_{0})(A\mid S)}\delta_{f}(S^{\prime},A,S)\bigg{]}$
and
[…]
$=$ $\mathrm{E}_{S\sim\omega(P_{0})(S;\pi^{*}(P_{0})),A\sim\pi^{*}(P_{0})}\delta_{f}(S^{\prime},A,S)-\mathrm{E}_{S\sim\omega(P_{\epsilon})(S;\pi^{*}(P_{\epsilon})),A\sim\pi^{*}(P_{\epsilon})}\bigg{[}\frac{\pi^{*}(P_{0})(A\mid S)}{\pi^{*}(P_{\epsilon})(A\mid S)}\delta_{f}(S^{\prime},A,S)\bigg{]}$
Here we apply $f(s)=V\big{(}s;\pi^{*}(P_{\epsilon}))$ in (16) and the equation can be rewritten as
$\Big{\langle}\omega(P_{\epsilon})\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}S;\pi^{*}(P_{0})\big{)},\,\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}\Big{\rangle}_{P_{0}}$
$=$ $\mathrm{E}_{P_{\epsilon}}V\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\mathrm{E}_{S\sim\omega(P_{0})(S;\pi^{*}(P_{0})),A\sim\pi^{*}(P_{0})}\bigg{[}\frac{\pi^{*}(P_{\epsilon})(A\mid S)}{\pi^{*}(P_{0})(A\mid S)}\Big{(}R+\gamma V\big{(}S^{\prime};\pi^{*}(P_{\epsilon}))-V\big{(}S;\pi^{*}(P_{\epsilon}))\Big{)}\bigg{]}$
$=$ $\mathrm{E}_{P_{\epsilon}}V\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\mathrm{E}_{P_{0},A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{\epsilon})\big{)}=\mathrm{E}_{P_{\epsilon}}V\big{(}S;\pi^{*}(P_{\epsilon})\big{)},$
in which
$\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}$ $=\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\big{[}Q\big{(}A,S;\pi^{*}(P_{0})\big{)}-V\big{(}S;\pi^{*}(P_{0})\big{)}\big{]}$
$=Q\big{(}A,S;\pi^{*}(P_{0})\big{)}\Big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)\Big{)}.$
Let us consider using the trick in the proof of Lemma A.3 again, the following ratio has both non-zero lower and upper bounds as
$\frac{\big{|}\omega(P_{\epsilon})\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}S;\pi^{*}(P_{0})\big{)}\big{|}}{\big{|}\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{|}}\leq\frac{1}{\underline{c}_{R}\operatorname{essinf}_{s\in\mathcal{S}}\inf_{a\in\mathcal{A}}\big{|}\pi^{*}(P_{\epsilon})(a\mid s)-\pi^{*}(P_{0})(a\mid s)\big{|}}=\frac{1}{\delta\underline{c}_{R}}$
and
$\frac{\big{\|}\omega(P_{\epsilon})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},1}}{\big{|}\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}\big{|}}\geq\frac{1}{\overline{c}_{R}\overline{c}_{\pi}}\frac{\delta\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{\pi}^{1/2}+\overline{c}_{\pi}^{-1}\delta}\geq\frac{\delta\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}.$
Thus, applying the reverse Cauchy inequality again, we obtain
$\big{\|}\omega(P_{\epsilon})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},2}^{2}\big{\|}\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},2}^{2}$ (18)
$\leq$ $\frac{4\underline{c}_{R}\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}{\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}\Bigg{(}\frac{\delta\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}+\frac{1}{\delta\underline{c}_{R}}\Bigg{)}^{2}$
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\Big{\langle}\omega(P_{\epsilon})\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}S;\pi^{*}(P_{0})\big{)},\,\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}\Big{\rangle}_{P_{0}}^{2}.$
Now by noticing that $\big{\|}\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},2}^{2}\geq\delta^{2}\underline{c}_{R}^{2}>0$. Let us show the inner product above will be $o_{P_{0}}(\epsilon)$. Indeed, by using the equations in the proof of Lemma A.3 and Lemma A.5, we obtain
$\Big{\langle}\omega(P_{\epsilon})\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}S;\pi^{*}(P_{0})\big{)},\,\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}\Big{\rangle}_{P_{0}}$
$=$ $\bigg{\langle}\omega(P_{\epsilon})\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}S;\pi^{*}(P_{0})\big{)},\,Q\big{(}A,S;\pi^{*}(P_{0})\big{)}\Big{(}\pi^{*}(P_{\epsilon})(A\mid S)-\pi^{*}(P_{0})(A\mid S)\Big{)}\bigg{\rangle}_{P_{0}}$
$=$ $\frac{1}{(1-\gamma)^{2}}\mathrm{E}_{P_{0}}\big{[}V\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-V\big{(}S;\pi^{*}(P_{0})\big{)}\big{]}$
$~{}~{}+\frac{1}{1-\gamma}\mathrm{E}_{P_{0},S\sim\omega(\boldsymbol{\cdot};\pi^{*}(P_{0}))}\bigg{[}\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\Big{[}Q\big{(}A,S;\pi^{*}(P_{0})\big{)}-Q\big{(}A,S;\pi^{*}(P_{\epsilon})\big{)}\Big{]}+V\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-V\big{(}S;\pi^{*}(P_{0})\big{)}\bigg{]}$
Here we omit the detailed steps in the last equation above, one can easily verify them as they are just equations. We have show that
$\big{\|}Q(P_{\epsilon})\big{(}a,s;\pi^{*}(P_{\epsilon})\big{)}-Q(P_{0})\big{(}a,s;\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},\infty}=o_{P_{0}}(\epsilon)$
in (12). Now, take advantage of our submodel in (13), under Assumption 3, we have
$\big{\|}V\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-V\big{(}S;\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},\infty}$ $=\Big{\|}\sup_{\pi\in\mathcal{P}}V(P_{\epsilon})(S;\pi)-\sup_{\pi\in\mathcal{P}}V(P_{0})(S;\pi)\Big{\|}_{P_{0},\infty}$
$\leq\sup_{\pi\in\mathcal{P}}\big{\|}V(P_{\epsilon})(S;\pi)-V(P_{0})(S;\pi)\big{\|}_{P_{0},\infty}$
$=\sup_{\pi\in\mathcal{P}}\big{\|}\mathrm{E}_{A\sim\pi}\omega(S;\pi)\big{(}1+\epsilon h_{R}(R\mid A,S)\big{)}\mathrm{d}P_{R,0}-\mathrm{E}_{A\sim\pi}\omega(S;\pi)\mathrm{d}P_{R,0}\big{\|}_{P_{0},\infty}$
$=\sup_{\pi\in\mathcal{P}}\big{\|}\epsilon\mathrm{E}_{S\sim P_{0},A\sim\pi}h_{R}(R\mid A,S)\big{)}\mathrm{d}P_{R,0}\big{\|}_{P_{0},\infty}=o_{P_{0}}(\epsilon).$
Therefore, we can furthermore refine (18) as
$\big{\|}\omega(P_{\epsilon})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},2}^{2}$
$\leq$ $\frac{4\underline{c}_{R}\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}{\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}\Bigg{(}\frac{\delta\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}+\frac{1}{\delta\underline{c}_{R}}\Bigg{)}^{2}$
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{\Big{\langle}\omega(P_{\epsilon})\big{(}S;\pi^{*}(P_{\epsilon})\big{)}-\omega(P_{0})\big{(}S;\pi^{*}(P_{0})\big{)},\,\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,S;\pi^{*}(P_{0})\big{)}\Big{\rangle}_{P_{0}}^{2}}{\big{\|}\mathrm{E}_{A\sim\pi^{*}(P_{\epsilon})}\mathbb{A}\big{(}A,\boldsymbol{\cdot};\pi^{*}(P_{0})\big{)}\big{\|}_{P_{0},2}^{2}}$
$\leq$ $\frac{4\underline{c}_{R}\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}{\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}\Bigg{(}\frac{\delta\gamma\sqrt{2\overline{c}_{\pi}^{-1}\underline{c}_{\pi}}\operatorname{BC}(S_{0},U(\mathcal{S}))}{\overline{c}_{R}\overline{c}_{\pi}\big{(}1+\overline{c}_{\pi}^{1/2}\big{)}}+\frac{1}{\delta\underline{c}_{R}}\Bigg{)}^{2}\frac{o_{P_{0}}(\epsilon^{2})}{\delta^{2}\underline{c}_{R}^{2}}=o_{P_{0}}(\epsilon^{2}),$
which leads to a contradiction, since we have (15) holds. Thus, we must have
$\big{|}\pi^{*}(P_{\epsilon})(a\mid s)-\pi^{*}(P_{0})(a\mid s)\big{|}=o_{P_{0}}(\epsilon)\qquad\text{ for all }\quad(a,s)\in\mathcal{A}\times\mathcal{S}_{\star}\quad\text{ with }\quad\mu(\mathcal{S}_{\star})>0$ (19)
if Assumption 3 holds. Simiarly, once apply $f(s)=V\big{(}s;\pi^{*}(P_{0}))$ in (17) and obtain the same conclusion above.
A.4 Proof of Theorem 3.1
Note that both $\mathcal{S}$ and $\mathcal{A}$ are compact. Thus, under the uniquess assumption in Theorem 3.1, (19) direclty implies $\mathbbl{\Delta}_{1}(\epsilon;\pi)=\mathbbl{\Delta}_{2}(\epsilon;\pi)=\mathbbl{\Delta}_{3}(\epsilon;\pi)=o_{P_{0}}(\epsilon)$ if we fix $\pi=\pi^{*}(P_{0})$. Therefore, using the decomposition in (11), we have
$\lim_{\epsilon\rightarrow 0}\frac{\Psi^{*}(P_{\epsilon})-\Psi^{*}(P_{0})}{\epsilon}$ $=\lim_{\epsilon\rightarrow 0}\left.\frac{\Psi(P_{\epsilon};\pi)-\Psi(P_{0};\pi)}{\epsilon}\right|_{\pi=\pi^{*}(P_{0})}+\lim_{\epsilon\rightarrow 0}\frac{o_{P_{0}}(\epsilon)}{\epsilon}$
$=\lim_{\epsilon\rightarrow 0}\left.\frac{\Psi(P_{\epsilon};\pi)-\Psi(P_{0};\pi)}{\epsilon}\right|_{\pi=\pi^{*}(P_{0})}.$
The above display guarantees that for any differentiable path $\{P_{\epsilon}:\epsilon\in\mathbb{R}\}$, the score function for $\Psi^{*}(P)$ is the exactly same as the score function for $\Psi(P;\pi)$ evaluated at $\pi=\pi^{*}(P_{0})$. Thus, the tangent space of $\Psi^{*}(P)$, denoted by $\mathcal{T}_{\Psi^{*}}$, will be the exactly same as the set of all elements in the tangent space of $\Psi(P;\pi)$ evaluated at $\pi=\pi^{*}(P_{0})$. It remains to show that $S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}$ minimizes the variances within $\mathcal{T}_{\Psi^{*}}$. Indeed, take any influence function in $\mathcal{T}_{\Psi^{*}}$, say $\operatorname{IF}\{\Psi(P^{*})\}\big{|}_{P=P_{0}}$, we have $\operatorname{IF}\{\Psi(P^{*})\}\big{|}_{P=P_{0}}=\operatorname{IF}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}$. Therefore,
$\mathrm{E}_{P_{0}}\big{[}\operatorname{IF}\{\Psi(P^{*})\}\big{|}_{P=P_{0}}\big{]}^{2}$ $=\mathrm{E}_{P_{0}}\big{[}\operatorname{IF}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}\big{]}^{2}\geq\mathrm{E}_{P_{0}}\big{[}S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}\big{]}^{2}.$
Since $S^{\text{eff, nonpar}}\{\Psi(P;\pi)\}\big{|}_{P=P_{0},\pi=\pi^{*}(P_{0})}$ is also a valid influence function for $\Psi^{*}(P)$ in $\mathcal{T}_{\Psi^{*}}$, thus it must be the efficient influence function.
A.5 Proof of Theorem 3.2
If Assumption 3.1 holds, we consider the following submodel in (13) as
$h_{R}(r\mid a\sim\pi,s)$ $=\mathds{1}\big{\{}\pi=\pi^{*},V(P_{0})(s;\pi)>Q(P_{0})(a,s;\pi)\big{\}}\big{(}1\wedge\operatorname{OR}(a,s)\big{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathds{1}\big{\{}\pi=\pi^{\star},V(P_{0})(s;\pi)\leq Q(P_{0})(a,s;\pi)\big{\}}\big{(}1\wedge\operatorname{OR}^{-1}(a,s)\big{)}.$
where the odd ratio is defined as
$\operatorname{OR}(a,s):=\frac{\mathrm{P}_{P_{0}}\big{(}V(P_{0})(s;\pi^{\star})\leq Q(P_{0})(a,s;\pi^{\star})\big{)}}{\mathrm{P}_{P_{0}}\big{(}V(P_{0})(s;\pi^{*})>Q(P_{0})(a,s;\pi^{*})\big{)}}.$
Since $\mathrm{E}_{P_{0}}[h_{R}(R\mid a\sim\pi,s)]=0$ for any policy $\pi\in\mathcal{P}$, it is a valid nontrivial submodel. However, in this submodel, take $\epsilon\downarrow 0$, the optimal policy is chosen as $\pi^{*}(P_{\epsilon})$ on $P_{\epsilon}$; while take $\epsilon\uparrow 0$, the optimal policy is chosen as $\pi^{\star}(P_{\epsilon})$ on $P_{\epsilon}$. Therefore, the expectation of the total variation distance between $\arg\max\pi(P_{\epsilon})$ and $\arg\max\pi(P_{0})$ will be a positive constant by Assumption 3.1, and the lower bound in Lemma A.3 implies we have $|\Psi^{*}(P_{\epsilon})-\Psi^{*}(P_{0})|=\frac{c_{0}}{\epsilon}$ for some stricly positive constant $c_{0}$, which implies $\Psi^{*}$ is not pathwise differentiable at $P_{0}$.
[…]
Auxiliary Lemmata
Lemma A.7 (Performance Difference Lemma, see Kakade & Langford (2002)).
Suppose that Assumption 2.1 and 2.1 hold. Then
$V(s_{0};\pi_{2})-V(s_{0};\pi_{1})=\frac{1}{1-\gamma}\mathrm{E}_{S\sim\omega(s_{0};\pi_{2})}\big{[}\mathrm{E}_{A\sim\pi_{2}(\cdot\mid S)}\mathbb{A}(A,S;\pi_{1})\big{]}.$
Lemma A.8 (Policy Decomposition Lemma, see Lemma 2 in Achiam et al. (2017)).
Suppose that Assumption 2.1 and 2.1 hold. For any function $f:\mathcal{S}\rightarrow\mathbb{R}$ and any policies $\pi_{1}$ and $\pi_{2}$, define
[…]
Then
[…]